Can you please help me with these problems, I am studying for my algebra exam and there are few things that i dont know how to do. If you could help it means a world to me. Thank you so much!
A=1/2bh Solve for b
F=5/9C+32 Solve for C
Can you please help me with Graphing Peace wise functions too.
1/2x+3/2 if x<1
-x+3 if x >1

I know its impossible to graph in here so it will be very nice if you just tell me how you solved it, because i kind of know how to graph it,

Thank you so much!

Answer by edjones(8007)

A=1/2bh Solve for b This is written incorrectly. It should be:
A=1/2 * bh
= bh/2
A=b * h/2
b=2A/h
.
Ed